# A Flat Minima Perspective on Understanding Augmentations and Model Robustness

### v1

@@ Abstract
Model robustness indicates a model’s capability to generalize well on unforeseen distributional shifts, including data corruption, adversarial attacks, and domain shifts. Data augmentation is one of the prevalent and effective ways to enhance robustness. Despite the great success of augmentations in different fields, a general theoretical understanding of their efficacy in improving model robustness is lacking. We offer a unified theoretical framework to clarify how augmentations can enhance model robustness through the lens of loss surface flatness and PAC generalization bound. Our work diverges from prior studies in that our analysis i) broadly encompasses much of the existing augmentation methods, and ii) is not limited to specific types of distribution shifts like adversarial attacks. We confirm our theories through simulations on the existing common corruption and adversarial robustness benchmarks based on the CIFAR and ImageNet datasets, as well as domain generalization benchmarks including PACS and OfficeHome.
1 Introduction
@@
Model robustness, which is a critical factor of deep models in applications requiring high reliability such as autonomous vehicles and medical diagnosis, entails maintaining performance against data distribution shifts. In the past decade, data augmentation has been widely used as a popular and pragmatic technique to enhance the model performance, as well as the robustness against data corruption, adversarial attacks, or even domain shifts augment_survey ; ME-ADA ; Mixup:2018 ; DeepAug10:2021 ; AutoAugment . The intuition of its efficacy relies on the belief that augmentations enrich the training data distribution, which allows models to easily extrapolate to unseen data distributions, which is the so-called generalization capability.
Figure 1: A conceptual overview of how augmentations relate to model robustness. Theorem 1 shows that the neighboring region of an input datapoint can be mapped to an equivalent region in the parameter space (and vice versa). Theorem 2 demonstrates that ensuring nonzero probability for an augmentation within this neighborhood can promote flatter minima in parameter space by effectively regulating the proximal regions of both the input and parameter spaces. Finally, Theorem 3 presents a generalization bound for models trained with augmentations, asserting that augmentations with dense representations near the original image foster a tighter generalization gap.
@@
Despite the utility of augmentations, a unified theory formally clarifying how augmentations enhance model robustness against diverse distribution shifts has not been well established. Prior analyses are mostly limited to either specific augmentations or adversarial robustness ME-ADA ; ADA ; AugmentNAdvGood ; NoLabelNAdvGood1 ; NoLabelNAdvGood2 ; aug_analysis_adv1 . Although some studies step forward from the aforementioned works, their analyses rely solely on empirical observations without providing a theoretical rationale, or they focus exclusively on certain categories of distribution shifts aug_analysis_cc1 ; aug_analysis_custom1 ; aug_analysis_dg . For instance, zhang2024duality links adversarial training with sharpness-aware minimization, but the results are confined to adversarial robustness.
@@
In this paper, we first present a set of theoretical explanations revealing how data augmentations can reinforce model robustness across general distributional changes in the perspective of flat minima. Our analysis has two main branches: i) First, we formally bridge the general form of augmentations to the improved generalization bound. To give a brief sketch of our development, we start to demonstrate the equivalence between the input space region covered by the augmented samples and the corresponding parameter space region with the same loss values (formalized by Theorem 1). Based on the equivalence, we then claim that the augmentations flatten the loss surface on the parameter space (Theorem 2), finally reaching to the improved generalization bound against distribution shifts via leveraging the flattened loss surface (Theorem 3). ii) Next, we evaluate existing augmentation methods across a range of robustness benchmarks, encompassing data distribution shifts caused by common corruptions, adversarial attacks, and domain changes. Our findings show that augmentations with sample coverage near the original image—which aligns with our assumption (Assumption 1)—tend to consistently enhance overall model robustness. In contrast, augmentations that rarely improve robustness under diverging distribution shifts exhibit vacant density in close proximity to the original image.
2 Preliminaries
2.1 Basic Notations
@@
$\mathcal{E}_{\mathcal{D}}(\theta):=\mathbb{E}_{(x,y)\sim\mathcal{D}}\big{[}\mathcal{L}(f(x;\theta),y)\big{]},\quad\hat{\mathcal{E}}_{\mathcal{D}}(\theta):=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f(x_{i};\theta),y_{i})$ (1)
respectively, where $\mathcal{E}_{\mathcal{D}}(\theta)$ and ^$\hat{\mathcal{E}}_{\mathcal{D}}(\theta)$ denote the true and the empirical risk.
2.1.1 Data Augmentation
Data augmentation $\mathcal{A}(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ augments a given input $x$ to augmented input $\tilde{x}:=\mathcal{A}(x)$. Let us further represent it with the difference between the original input and the augmented one: $\tilde{x}:=\mathcal{A}(x)=x+\delta$, where $\delta\in\mathbb{R}^{n}$.  [^1] By formulating augmentation as an additive perturbation, we transform the perturbation on the parameter space to the one on the input space. Details are in the following section. Also, we define $\mathcal{P}_{\mathcal{A}}(\tilde{x}|x)$ as the probability density function of an augmented sample $\tilde{x}$ given $x$. Finally, we let $\tilde{\mathcal{D}}:=\{(\mathcal{A}(x),y):(x,y)\sim\mathcal{D}\}$ represent the augmented dataset.
2.2 Model Flatness
2.2.1 Definition
@@
where $\Delta\in\mathbb{R}^{p}$ perturbs model parameter $\theta$ within radius $\gamma>0$ to maximize the loss.
2.2.2 Sharpness-aware Minimization
The most popular principal way for finding flat minima is utilizing sharpness-aware minimization (SAM) optimizer SAM:2021 , which formally transforms loss minimization into a min-max optimization:
@@ 2.2.2 Sharpness-aware Minimization
As formulated in the work of SAM:2021 , when adding the loss value at $\theta$, only the first term of (2) remains, yielding the objective function in (3). Through the minimization of the maximization term, SAM aims to find the minima $\theta$ with flatter loss surfaces around $\gamma$ radius.
3 How Can Augmentations Enhance Robustness?
In this section, we provide a rigorous link between data augmentations and improved generalization capability. Our main claims are twofold: i) Translation between equivalent input manifold and parameter manifold, ii) Association of flatter loss in the input space via augmentations with flatness in the parameter space, leading to a reduced generalization gap.
3.1 Translation Between Input and Parameter Perturbations
@@
Conversely, the same intuition in reverse applies for parameter perturbations. Consider a perturbation $\Delta\in\mathbb{R}^{p}$ around $\theta$. For any $\|\Delta\|\leq\gamma$, we posit an analogous compensatory region $\mathcal{C}_{\mathcal{X}}$ in the input space such that $f(x;\theta+\Delta)=f(x+\delta;\theta)$. Specifically, there is some input shift $\delta\in\mathcal{C}_{\mathcal{X}}$ that replicates the same output change induced by $\Delta$.
Our goal is to formalize this duality of perturbations in both the input and parameter spaces. We now present the notion of functional compensatory sets, which identifies the precise sets of input shifts needed to compensate for every parameter perturbation of norm up to $\gamma$, and vise versa.
3.1.1 Functional Compensatory Sets
@@
This set is not necessarily unique, and it characterizes all possible input shifts that replicate the effect of parameter perturbations within norm $\gamma$.
@@
3.1.2 Formalizing the Input–Parameter Perturbation Duality
@@
Let $\mathbf{J}$ be the Jacobian of a function $f:\mathbb{R}^{n}\times\mathbb{R}^{p}\rightarrow\mathbb{R}^{c}$, where $\mathbb{R}^{n},\mathbb{R}^{p},$ and $\mathbb{R}^{c}$ are the input, parameter, and output spaces. The Jacobian matrix $\mathbf{J}\in\mathbb{R}^{c\times(n+p)}$ can be decomposed into its input side and the parameter side partial derivatives, i.e. $\mathbf{J}=\left[\mathbf{J_{x}}\ \ \mathbf{J}_{\boldsymbol{\theta}}\right]$, where $\mathbf{J_{x}}\in\mathbb{R}^{c\times n}$ and $\mathbf{J}_{\boldsymbol{\theta}}\in\mathbb{R}^{c\times p}$. We will use the notation $\mathbf{J_{x}}(x_{i})$ and $\mathbf{J}_{\boldsymbol{\theta}}(\theta)$ when we need to explicitly represent the evaluation of $\mathbf{J_{x}}$ and $\mathbf{J}_{\boldsymbol{\theta}}$ at point $x_{i}\in\mathbb{R}^{n}$ and $\theta\in\mathbb{R}^{p}$, respectively. Finally, let $\sigma_{\min}^{x_{i}}$ and $\sigma_{\max}^{\theta}$ represent the mininum and the maximum values of the matrix $\mathbf{J_{x}}(x_{i})$ and $\mathbf{J}_{\boldsymbol{\theta}}(\theta)$. Leveraging Sard’s theorem Sard together with Taylor expansion, we construct two sets—$\mathcal{S}_{\mathcal{X}}$ and $\mathcal{S}_{\Theta}$—that fulfill the definition of functional compensatory sets $\mathcal{C}_{\mathcal{X}}$ and $\mathcal{C}_{\Theta}$ as follows:
@@
3.2 Linking Augmentations to Model Flatness and Generalization
Built upon the theorems above, we now formalize how can augmentations lead to flatter loss landscape around the minima and the improved generalization capability. Before providing the details, let us rephrase the formal definition of flat minima, which is called $b$-flat minima shi2021overcoming :
@@ 3.2 Linking Augmentations to Model Flatness and Generalization
For input $x$, an augmentation $\mathcal{A}$ induces a distribution of variants $\tilde{x}$ near $x$. Let $\gamma_{\mathcal{A}}$ be the largest radius of a ball around $x$ with non-zero probability. If $\gamma_{\mathcal{A}}>0$, we say $\mathcal{A}$ covers the local neighborhood. We assume this property for augmentations currently considered and show in Section 4 that it is critical for the robustness of augmentations.
Assumption 1.
@@
Let $\tilde{\mathcal{D}}$ represent the augmented dataset with respect to $\mathcal{D}$ and $\gamma_{\mathcal{A}}$ the radius value satisfying Assumption 1. Now, let $\Theta^{*}$ and $\tilde{\Theta}^{*}$ be the sets of the optimal model parameters whose elements $\theta^{*}\in\Theta^{*}$ and $\tilde{\theta}^{*}\in\tilde{\Theta}^{*}$ satisfy the following equalities, $\mathcal{E}_{\mathcal{D}}(\theta^{*})=0$ and $\mathcal{E}_{\mathcal{\tilde{D}}}(\tilde{\theta}^{*})=0$, respectively. Then our claim is that the minimum $b$-flatness among the solution parameters in $\tilde{\Theta}^{*}$, shows large $b$ (flatter) than the minimum $b$-flatness among the solutions in $\Theta^{*}$, which is trained on $\mathcal{D}$:
@@
The final linkage to the generalization capability is straightforward by relying on the prior theoretical results that bridge Robust Risk Minimization (RRM) and the generalization bound SWAD:2021 . RRM aims to find a flat minimum, where its surrounding region shows flat loss surface. Based on Theorem 2, augmentations make the optimal parameter locate at a flatter surface, and it implies the improved generalization bound (referred to Theorem 3). Let us define $\gamma_{\Theta}:=\left(\max_{x\in\mathcal{D}}\sigma_{\max}^{\theta}/\sigma_{\min}^{x}\right)^{-1}\cdot\gamma_{\mathcal{A}}$, then the following theorem holds:
@@
(Generalization bound) Given $M$ covering sets $\{\Theta_{k}\}_{k=1}^{M}$ of parameter space $\Theta$ with $\Theta=\bigcup_{k=1}^{M}\Theta_{k}$ and $\text{diam}(\Theta)=\sup_{\theta,\theta^{\prime}\in\Theta}||\theta-\theta^{\prime}||_{2}$, where $M=\left\lceil\frac{\text{diam}(\Theta)}{\gamma_{\Theta}}\right\rceil^{d}$, and VC dimension $v_{k}$ for each $\Theta_{k}$, the following inequality holds with probability at least $1-\delta$:
$\mathcal{E}_{\mathcal{T}}(\tilde{\theta}^{*})$ <^$<\hat{\mathcal{E}}_{\mathcal{\tilde{D}}}(\tilde{\theta}^{*})+\frac{1}{2}\textbf{Div}(\mathcal{D},\mathcal{T})+\max_{k\in[1,M]}\left[\sqrt{\frac{v_{k}\log(N/v_{k})}{N}+\frac{\log(M/\delta)}{N}}\right],$
@@
(Augmentations improve generalization against data distribution shifts) The theorem implies that the minimization of empirical loss for the augmented dataset, i.e., ^$\hat{\mathcal{E}}_{\mathcal{\tilde{D}}}(\tilde{\theta}^{*})$, directly aims the tighter generalization bound on target distribution $\mathcal{T}$. Also, the term $M$ is related to $\gamma_{\mathcal{A}}$, which measures how augmented data distribution covers a wider range (referred to Assumption 1). When augmentation covers a wider range around the original sample, i.e., a larger $\gamma_{\mathcal{A}}$, it suppresses $M$, leading to the smaller last term of generalization bound. These interpretation elucidates how augmentations enhance the generalization capability against unseen data distribution shifts.
4 Empirical Examination on Proximal Density, Model Flatness, and Generalization
In this section, we put our theory to test by examining how augmentations influence model robustness. Earlier, we established that any augmentation assigning non-zero probability to samples lying in a neighborhood around the original image (Assumption 1) maps to an analogous region in parameter space (Theorem 1), which in turn induces flatter minima (Theorem 2) and yields a tighter generalization bound (Theorem 3). Therefore, augmentations with rich proximal density should yield strong robustness gains under diverse distribution shifts, whereas augmentations that fail to enhance robustness should suffer from near-zero representations in the proximal region of the original image.
4.1 Augmentation Methods to Be Considered
Following the taxonomy of augment_survey , augmentations can be categorized into model-free, model-based, and policy-based augmentations. Among the immense existing augmentation methods, we focus on well-known or advanced methods rather than basic or elementary augmentations. For the model-free augmentation type, we select AugMix AugMix9:2021 and PixMix PixMix:2022 , which are advanced forms of mixing multiple clean images to the original image. For the model-based augmentation type, we choose StyleAug StyleAug and DeepAug DeepAug10:2021 , which utilize image-to-image models to diversify the style of the clean images. Finally, among the policy-based augmentation type, we consider RandAugment RandAugment , which learns the policy for augmenting images.
4.2 Tests on Proximal Density
We first measure each augmentation’s proximity (or “proximal density”) to the original image via observing the empirical cumulative distribution function (eCDF) of $L_{2}$ distances. Specifically, we compute how quickly the eCDF accumulates at smaller values, which provides a clear indicator of each augmentation’s density near the original image. Formally, if $\{d_{i}\}_{i=1}^{N}$ are these distances, then the eCDF is defined as $F_{N}(\gamma)=\tfrac{1}{N}\sum_{i=1}^{N}\mathbf{1}\{d_{i}\leq\gamma\}$. High values in the lower-distance region of $F_{N}(\gamma)$ signifies that a large fraction of augmented samples remain close to the original images. Conversely, an augmentation whose representations are spread out will exhibit small values to small gamma, indicating sparser coverage.
As illustrated in Figure 2 and Table 1, on CIFAR-100 dataset AugMix have dense representations near the original, RandAug and PixMix show sparse but non-zero clustering, and DeepAug and StyleAug exhibit nearly zero coverage at low distances. On the ImageNet dataset, the results are analogous to CIFAR, except that DeepAug’s eCDF now accumulates much more rapidly and has dense representations in the proximal region unlike CIFAR.
Figure 2: Distance histograms on CIFAR-100 illustrating augmentation density (top: entire distribution, bottom: near origin). From the highest to the lowest coverage near the origin: AugMix is dense, RandAug and PixMix are sparse, and DeepAug and StyleAug have zero. A dense concentration of histogram close to the original suggests a higher likelihood that $\gamma_{\mathcal{A}}\gg 0$, whereas an absence of samples in the low-distance region indicates $\gamma_{\mathcal{A}}=0$ (referring to Assumption 1).
Table 1: Empirical CDF $\big{(}F_{N}(\gamma)\big{)}$ for selected distance thresholds $(\gamma)$, evaluated across various augmentation methods on CIFAR-10/100 and ImageNet. Red entries indicate zero or negligible probability $(<5\cdot 10^{-5})$, and green entries exceed 0.01, illustrating how densely each augmentation stays near the original image.
Distance CIFAR-100 Distance ImageNet AugMix RandAug PixMix DeepAug StyleAug AugMix RandAug PixMix DeepAug StyleAug $\gamma=0.01$ 0.0024 0.0060 0.0011 0.0000 0.0000 $\gamma=20.0$ 0.0005 0.0001 0.0002 0.0005 0.0000 $\gamma=0.05$ 0.0107 0.0060 0.0011 0.0000 0.0000 $\gamma=30.0$ 0.0031 0.0009 0.0013 0.0028 0.0000 $\gamma=0.10$ 0.0202 0.0061 0.0012 0.0000 0.0000 $\gamma=40.0$ 0.0113 0.0034 0.0054 0.0110 0.0001 $\gamma=0.50$ 0.0993 0.0378 0.0012 0.0001 0.0000 $\gamma=50.0$ 0.0317 0.0104 0.0169 0.0311 0.0008
4.3 Tests on Model Flatness
We now explore how an augmentation’s density near the original image influences flatness in the learned model parameters, linking our results from Theorems 1 and 2 with empirical flatness.
Flatness metrics. We assess each model using three quantitative measures: PAC-Bayesian flatness measure $\mu_{\text{PAC-bayes}}$ PACBayes , local sharpness $\epsilon_{\text{sharp}}$ eps-sharpness , and Loss-Pass Filter (LPF) metric, a strong indicator of generalization LPF . Each of these metrics reflects the difference between the “proximal loss” around a model parameter and its original loss, which aligns well with our theoretical framework. We therefore avoid purely “pinpoint-based” measures (e.g., maximum Hessian eigenvalues or their traces), as they do not fully capture our theorems’ emphasis on local neighborhoods in parameter space. For more comprehensive descriptions of each metric, refer to Appendix C.1.1.
Experimental results. We train WideResNet-40-2 architecture on CIFAR-10/100 datasets adopting the five augmentations. For further details on training, see Appendix C.1.1. Table 2 presents each model’s flatness on CIFAR, under different augmentation settings. In line with our theorems and assumption, methods that retain nonzero probability of producing samples close to the original image (AugMix, RandAug, PixMix) yield flatter minima than ERM, irrespective of whether their coverage is dense or sparse. Meanwhile, augmentations that fail to exceed ERM in achieving flatter minima (DeepAug, StyleAug) are those exhibiting near-zero proximal density. These observations align with Theorems 1, 2, and Assumption 1, underscoring that maintaining a nontrivial concentration of augmented samples near the original input is a critical aspect for finding flatter minima.
In the (tiny)ImageNet setting, augmentations with high proximal density (Table 1) again produce substantially flatter minima than plain ERM. All methods except StyleAug achieve this benefit, and DeepAug does so as well in this case. Detailed table statistics are at Appendix C.1.
Table 2: Flatness metrics of ERM and different augmentations on CIFAR benchmarks (:absent:$\downarrow$:↓ : The lower the better, i.e., flatter minimum). Overall, augmentations with nonzero probability near the original image consistently yield flatter minima, whereas those with near-zero coverage struggle to outperform ERM on all the flatness measures.

### v3

@@ Abstract
Model robustness indicates a model’s capability to generalize well on unforeseen distributional shifts, including data corruptions and adversarial attacks. Data augmentation is one of the most prevalent and effective ways to enhance robustness. Despite the great success of the diverse augmentations in different fields, a unified theoretical understanding of their efficacy in improving model robustness is lacking. We theoretically reveal a general condition for label-preserving augmentations to bring robustness to diverse distribution shifts through the lens of flat minima and generalization bound, which de facto turns out to be strongly correlated with robustness against different distribution shifts in practice. Unlike most earlier works, our theoretical framework accommodates all the label-preserving augmentations and is not limited to particular distribution shifts. We substantiate our theories through different simulations on the existing common corruption and adversarial robustness benchmarks based on the CIFAR and ImageNet datasets.
Codehttps://github.com/pyoo96/aug-flatmin-robustness
Introduction
@@
Model robustness, which is a critical factor of deep models in applications requiring high reliability, such as autonomous vehicles and medical diagnosis, entails maintaining performance against data distribution shifts. In the past decade, data augmentation has been widely used as a popular and pragmatic technique to enhance the model performance, as well as the robustness against data corruption, adversarial attacks, or even domain shifts (augment_survey; ME-ADA; PAC_DG1; DeepAug10:2021). The intuition of its efficacy relies on the belief that augmentations enrich the training data distribution, which allows models to easily extrapolate to unseen data distributions, which is the so-called generalization capability.
@@
Despite the utility of augmentations, a unified theory formally clarifying how augmentations can enhance model robustness against diverse distribution shifts has not been well established. Prior analyses are mostly confined to either particular augmentations or adversarial robustness (ME-ADA; AugmentNAdvGood; NoLabelNAdvGood1; aug_analysis_adv1). Although some studies step forward from the aforementioned works, their analyses rely solely on empirical observations without a theoretical rationale, or the focus is exclusively on certain categories of distribution shifts (zhang2024duality; aug_analysis_custom1; aug_analysis_dg).
@@
In this paper, we offer a series of theoretical insights, including a sufficient condition (termed the PSA condition), that explains how label-preserving data augmentations can bolster model robustness against general distributional shifts through the lens of flat minima. Our analysis has two main branches: i) First, we mathematically bridge the general form of label-preserving augmentations to the improved generalization bound. To give a brief sketch of our development, we start to demonstrate the equivalence between the input space region covered by the augmented samples and the corresponding parameter space region with the same loss values (formalized by Theorem 1). Based on the equivalence, we then claim that the augmentation satisfying the PSA condition flattens the loss surface on the parameter space (Theorem 2), finally reaching to the improved generalization bound against distribution shifts via leveraging the flattened loss surface (Theorem 3). ii) Next, we validate our theoretical findings by evaluating existing augmentation methods across different robustness benchmarks, encompassing data distribution shifts caused by common corruptions and adversarial attacks. Our findings show that when augmentations have non-negligible sample coverage near the original image–which aligns with the PSA condition–they consistently enhance model robustness. In contrast, when augmentations fail to improve robustness, they exhibit negligible density near the original image, indicating that PSA serves as both a sufficient condition for robustness and a factor highly correlated with robustness.
Figure 1: A conceptual overview of how augmentations can relate to model robustness. Theorem 1 shows that the neighboring region of an input datapoint can be mapped to an equivalent region in the parameter space (and vice versa). Theorem 2 demonstrates that ensuring a nonzero probability for an augmentation within this neighborhood (PSA condition) promotes flatter minima in parameter space by effectively regulating the proximal regions of both the input and parameter spaces. Finally, Theorem 3 presents a generalization bound for models trained with augmentations, insisting that augmentations with dense representations near the original image foster a tighter generalization gap.
Preliminaries
Basic Notations
@@
$\mathcal{E}_{\mathcal{D}}(\theta):=\mathbb{E}_{(x,y)\sim\mathcal{D}}\big[\mathcal{L}(f(x;\theta),y)\big]$ (1)
$\hat{\mathcal{E}}_{\mathcal{D}}(\theta):=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}(f(x_{i};\theta),y_{i})$ (2)
respectively, where $\mathcal{E}_{\mathcal{D}}(\theta)$ and $\hat{\mathcal{E}}_{\mathcal{D}}(\theta)$ denote the true and the empirical risk.
Data Augmentation
Data augmentation $\mathcal{A}(\cdot):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ augments a given input $x$ to augmented input $\tilde{x}:=\mathcal{A}(x)$. Let us further represent it with the difference between the original input and the augmented one: $\tilde{x}:=\mathcal{A}(x)=x+\delta$, where $\delta\in\mathbb{R}^{n}$.  [^1] By formulating augmentation as an additive perturbation, we transform the perturbation on the parameter space to one on the input space. Details are in the following section. Also, we define $\mathcal{P}_{\mathcal{A}}(\tilde{x}|x)$ as the probability density function of an augmented sample $\tilde{x}$ given $x$. Finally, we let $\tilde{\mathcal{D}}:=\{(\mathcal{A}(x),y):(x,y)\sim\mathcal{D}\}$ represent the augmented dataset.
Model Flatness
Definition
@@
where $\Delta\in\mathbb{R}^{p}$ is the perturbation around model parameter $\theta$ that maximally increases loss within a radius $\gamma>0$.
Sharpness-aware minimization
The most popular principal way for finding flatter minima is Sharpness-Aware Minimization (SAM) (SAM:2021), which formally transforms loss minimization into a min-max optimization:
@@
As formulated in the work of (SAM:2021), when adding the loss value at $\theta$, only the first term of (3) remains, yielding the objective function in (4). Through the minimization of the maximization term, SAM aims to find the minima $\theta$ with flatter loss surfaces around $\gamma$ radius.
How Can Augmentations Enhance Robustness?
In this section, we provide a rigorous link between data augmentations and improved generalization capability. Our main claims are twofold: i) Translation between equivalent input manifold and parameter manifold, ii) Association of flatter loss in the input space via augmentations with parameter space flatness, leading to a reduced generalization gap.
Translation Between Input and Parameter Perturbations
@@
Conversely, the same intuition in reverse applies to parameter perturbations. Consider a perturbation $\Delta\in\mathbb{R}^{p}$ around $\theta$. For any $\|\Delta\|\leq\gamma$, we posit an analogous compensatory region $\mathcal{C}_{\mathcal{X}}$ in the input space such that $f(x;\theta+\Delta)=f(x+\delta;\theta)$. Specifically, there is some input shift $\delta\in\mathcal{C}_{\mathcal{X}}$ that replicates the same output change induced by $\Delta$.
Our goal is to formalize this duality of perturbations in both the input and parameter spaces. We now present the notion of functional compensatory sets, which identifies the precise sets of input shifts needed to compensate for every parameter perturbation of norm up to $\gamma$, and vice versa.
Functional Compensatory Sets
@@
This set $\mathcal{C}_{\mathcal{X}}$ is not necessarily unique, and it characterizes all possible input shifts that replicate the effect of parameter perturbations within norm $\gamma$.
@@
Formalizing the Input–Parameter Perturbation Duality
@@
Let $\mathbf{J}$ be the Jacobian of a function $f:\mathbb{R}^{n}\times\mathbb{R}^{p}\rightarrow\mathbb{R}^{c}$, where $\mathbb{R}^{n},\mathbb{R}^{p},$ and $\mathbb{R}^{c}$ are the input, parameter, and output spaces. The Jacobian matrix $\mathbf{J}\in\mathbb{R}^{c\times(n+p)}$ can be decomposed into its input side and the parameter side partial derivatives, i.e. $\mathbf{J}=\left[\mathbf{J_{x}}\ \ \mathbf{J}_{\boldsymbol{\theta}}\right]$, where $\mathbf{J_{x}}\in\mathbb{R}^{c\times n}$ and $\mathbf{J}_{\boldsymbol{\theta}}\in\mathbb{R}^{c\times p}$. We will use the notation $\mathbf{J_{x}}(x_{i})$ and $\mathbf{J}_{\boldsymbol{\theta}}(\theta)$ when we need to explicitly represent the evaluation of $\mathbf{J_{x}}$ and $\mathbf{J}_{\boldsymbol{\theta}}$ at point $x_{i}\in\mathbb{R}^{n}$ and $\theta\in\mathbb{R}^{p}$, respectively. Finally, let $\sigma_{\min}^{x_{i}}$ and $\sigma_{\max}^{\theta}$ represent the mininum and the maximum values of the matrix $\mathbf{J_{x}}(x_{i})$ and $\mathbf{J}_{\boldsymbol{\theta}}(\theta)$.
[…]
@@
Distance CIFAR-10 CIFAR-100 AugMix RandAug PixMix DeepAug StyleAug AugMix RandAug PixMix DeepAug StyleAug $\gamma=0.01$ 0.0025 0.0064 0.0014 0.0000 0.0000 0.0024 0.0060 0.0011 0.0000 0.0000 $\gamma=0.05$ 0.0100 0.0064 0.0015 0.0000 0.0000 0.0107 0.0060 0.0011 0.0000 0.0000 $\gamma=0.10$ 0.0205 0.0065 0.0015 0.0000 0.0000 0.0202 0.0061 0.0012 0.0000 0.0000 $\gamma=0.50$ 0.0959 0.0391 0.0015 0.0001 0.0000 0.0993 0.0378 0.0012 0.0001 0.0000
Table 1: Empirical CDF $\big(F_{N}(\gamma)\big)$ for selected distance thresholds $(\gamma)$ on CIFAR datasets, illustrating sample density near the original image over different augmentation methods. The higher the eCDF value, the stronger an augmentation satisfies PSA condition. Sorted in descending order. (Italic entries have near-zero probability $(<5\cdot 10^{-5})$, and bold entries exceeds 0.01.)
Linking Augmentations to Model Flatness and Generalization
Built upon the theorems above, we now formalize how augmentations can lead to a flatter loss landscape around the minima and the improved generalization capability. Before providing the details, let us rephrase the formal definition of flat minima, which is called $b$-flat minima (shi2021overcoming):
@@
For input $x$, an augmentation $\mathcal{A}$ induces a distribution of variants $\tilde{x}$ near $x$. Let $\gamma_{\mathcal{A}}$ be the largest radius of a ball around $x$ with nonzero probability density. If $\gamma_{\mathcal{A}}>0$, we say $\mathcal{A}$ covers the local neighborhood. We now formalize this property and later show that it is critical for robustness.
Condition.
(Proximal-Support Augmentation (PSA))
@@
PSA is a condition that covers any label-preserving augmentation and ties it to flat minima. It serves as a testable rule by setting a local coverage requirement around $x$. Discarding dependencies on the model’s interpolation ability between data points, PSA is a strong condition for zero loss over all neighborhoods of inputs for any architectures.
Let $\gamma_{\mathcal{A}}$ be the radius value for PSA, and $\tilde{\mathcal{D}}$ represent the augmented dataset with respect to $\mathcal{D}$ under PSA condition. Let $\Theta^{*}$ and $\tilde{\Theta}^{*}$ be the sets of the optimal model parameters whose elements $\theta^{*}\in\Theta^{*}$ and $\tilde{\theta}^{*}\in\tilde{\Theta}^{*}$ satisfy the following equalities, $\mathcal{E}_{\mathcal{D}}(\theta^{*})=0$ and $\mathcal{E}_{\mathcal{\tilde{D}}}(\tilde{\theta}^{*})=0$, respectively. Then our claim is that the minimum $b$-flatness among the solution parameters in $\tilde{\Theta}^{*}$, shows large $b$ (flatter) than the minimum $b$-flatness among the solutions in $\Theta^{*}$, which is trained on $\mathcal{D}$:
@@
The final linkage to the generalization capability is straightforward by relying on the prior theoretical results that bridge Robust Risk Minimization (RRM) and the generalization bound (SWAD:2021). As RRM targets flat minima, Theorem 2 shows that augmentations further flatten the optimum’s neighborhood, and directly tightens the bound in Theorem 3. Let us define $\gamma_{\Theta}:=\left(\max_{x\in\mathcal{D}}\sigma_{\max}^{\theta}/\sigma_{\min}^{x}\right)^{-1}\cdot\gamma_{\mathcal{A}}$, then the following theorem holds:
@@
[…]
$+\max_{k\in[1,M]}\left[\sqrt{\frac{v_{k}\log(N/v_{k})}{N}+\frac{\log(M/\delta)}{N}}\right],$
@@
[…]